Question 13 of 14 (5 points) View problem in a pop-up 9.4 Section Exercise 18a,b,c,d,e The percentages of on-time arrivals fo Find the critical values in ascending order and round the answer to two decimal places
Answer #1

necessary calculations:-

n_{1}=233,\hat{p_{1}}=\frac{165}{233}=0.7082

n_{2}=312,\hat{p_{2}}=\frac{241}{312}=0.7724

\hat{p}=\frac{165+241}{233+312}=0.745

hypothesis:-

H_{0}:p_{1}-p_{2}=0

H_{1}:p_{1}-p_{2}\neq0

test statistic be:-

z=\frac{\hat{p_{1}}-\hat{p_{2}}}{\sqrt{\hat{p}(1-\hat{p})(\frac{1}{n_{1}}+\frac{1}{n_{2}})}}=\frac{0.7082-0.7724}{0.745(1-0.745*(\frac{1}{233}+\frac{1}{312})}=-1.70

**THE PART YOU NEEDED**

z critical value for alpha= 0.05, both tailed test , both tailed test be:-

\mathbf{z_{crit}=-1.96,1.96}

decision:-

\left | z \right |=1.70<\left | z_{calc} \right |=1.96

so, we fail to reject the null hypothesis and conclude that there is not sufficient evidence to claim that  there a difference in proportions

***in case of doubt, comment below. And if u liked the solution, please like.